given the sum of two numbers is thirty and that twice the first number is 3 more than the second,find the largest number
Answer by checkley77(12844)

Let x & y be the two numbers.
x+y=30 or x=30-y
2x=y+3
2(30-y)=y+3
60-2y=y+3
-2y-y=3-60
-3y=-57
y=-57/-3
y=19 ans for the larger number.
x+19=30
x=30-19
x=11 ans for the smaller number.
Proof:
2*11=19+3
22=22